A 50.0 g Super ball travelling at 25.0 m/s bounces off a brick wall and rebounds at 22.0 m/s. A high-speed camera records this event. If the ball is in contact with the wall for 3.50 s, what is the magnitude of the average acceleration of the ball during this time interval?
1
Expert's answer
2020-10-05T10:54:49-0400

Solution

Let us consider initially ball is moving in -x axis and after rebounce it moved in +x direction therefore velocities are "v_i=-25\widehat {i} m/s"

And final velocity "v_f=22\widehat{i} m/s"

Time interval "\Delta t=3.50 sec"


Therefore average acceleration

"\overrightarrow{a}=\frac{v_f-v_i}{\Delta t}=\frac{22-(-25)}{3.50}\\ \space\space\space\space=13.43 m/s^2."